# Simplified and Generalized Masked Diffusion for Discrete Data

### v2

@@ 8 Conclusion
In this work, we revisit masked diffusion models, focusing on a flexible continuous-time formulation. Existing works in this area are not easily accessible to non-specialists and present ELBOs that are difficult to optimize, often resulting in performance that is not competitive with continuous diffusions and autoregressive models. The framework we propose provides a very simple expression of the ELBO as a weighted integral of cross-entropy losses. Additionally, we propose a generalized masked diffusion formulation (GenMD4), where the masking schedule depends on the current state of the process, and derive its corresponding ELBO. On text data, our MD4 models outperform existing discrete and continuous diffusion models. For pixel-level image modeling, we significantly improve discrete diffusion results, outperforming similar-sized AR models and achieving comparable likelihoods to continuous diffusion models such as VDM. GenMD4 provides further improvements in terms of likelihoods over the state-independent case.
Although we have improved masked diffusion models, they still suffer from limitations. First, in some tasks such as text8, masked diffusions are not yet competitive with autoregressive models. We conjecture that this is because autoregressive models can better leverage model capacity since they only require learning one order. It would be interesting to develop better architectures for discrete diffusions. Moreover, GenMD4 is promising, but it can easily overfit to the dataset, making it less effective for zero-shot transfer compared to simpler versions. Additionally, inference with a state-dependent schedule is more challenging.
@@ H.3 Score parameterization
$\mathcal{L}_{\infty}=\int_{t(1)}^{1}\mathbb{E}_{q_{t|0}(k|x_{0})}\Big{[}\sum_{j\neq k}Q(t)_{jk}\Big{(}s_{\theta}(k)_{j}-\frac{q_{t|0}(j|x_{0})}{q_{t|0}(k|x_{0})}\log s_{\theta}(k)_{j}+\psi\Big{(}\frac{q_{t|0}(j|x_{0})}{q_{t|0}(k|x_{0})}\Big{)}\Big{)}\Big{]}\mathrm{d}t,$
@@
Note that only the result under the case $k=m$ is needed. This is because when $x_{t}$ is unmasked, at any time between 00 and $t$, the state must stay unchanged and remain $x_{0}$. As a result, $\mathrm{KL}({q(x_{t-\Delta t}|x_{t},x_{0})}\|{p_{\theta}(x_{t-\Delta t}|x_{t})})=0$ for $x_{t}\neq m$. From (15), we know $Q(t)_{jk}=\beta(t)(\delta_{mk}-\delta_{jk})$. Combining (24) and (H.3), we get
@@
This is the same as the conditional reverse transition rate used in Campbell et al. [48, Eq. (22)]—note that their time $t$ is reversed, and the rate matrix was therefore in the form $R^{x_{0}}(t)=\frac{1}{1-t}e_{m}(x_{0}-e_{m})^{\top}$.
@@ I.2 Training and gradient estimation
$\mathcal{L}_{\infty}=\int_{0}^{1}\Big{(}\frac{\alpha_{t}^{\prime}}{1-\alpha_{t}}\Big{)}^{\top}\mathbb{E}_{q(x_{t}|x_{0})}\left[\sum_{n:x_{t}^{(n)}=m}(x_{0}^{(n)}-\mu_{\theta}^{(n)}(x_{t},t)+x_{0}^{(n)}(x_{0}^{(n)})^{\top}\log\mu_{\theta}^{(n)}(x_{t},t))\right]\mathrm{d}t.$
@@
$\mathcal{L}_{\infty}=-\int_{0}^{1}\frac{1}{t}w^{\top}\mathbb{E}_{q(x_{t}|x_{0})}\left[\sum_{n:x_{t}^{(n)}=m}(x_{0}^{(n)}-\mu_{\theta}^{(n)}(x_{t},t)+x_{0}^{(n)}(x_{0}^{(n)})^{\top}\log\mu_{\theta}^{(n)}(x_{t},t))\right]\mathrm{d}t,$
@@
$-\frac{1}{2t}\left\{g(x_{t}^{1},x_{0})+g(x_{t}^{2},x_{0})+\left(\nabla_{w}\log q(x_{t}^{1}|x_{0})-\nabla_{w}\log q(x_{t}^{2}|x_{0})\right)\left[f(x_{t}^{1},x_{0})-f(x_{t}^{2},x_{0})\right]\right\}$
@@
J.3 Images
@@
We have observed that the cosine schedule leads to better sample quality so we used it to train a cheaper model for sample visualization. This model differs from the one that achieves best likelihood in that we used 8 residual blocks (in one side of the UNet structure) and a 20-layer diffusion transformer in the middle. All other configurations are kept the same.
@@
K.3 Additional unconditional generation from MD4 trained on ImageNet 64$\times$64
@@
K.4 Additional unconditional generation from MD4 trained on OpenWebText
@@
K.4.1 MD4-M unconditional sample 1: 1024 tokens
@@
K.4.2 MD4-M unconditional sample 2: 1024 tokens
@@
K.5 Conditional generation from MD4 trained on OpenWebText
@@
Figure 9: Conditionally generated text samples from MD4-M. Top: MD4-M trained with linear masking schedule; Bottom: MD4-M trained with cosine masking schedule. Context text shown in blue, model-generated text in black.
K.6 Effect of discretization on zero-shot perplexity

### v4

@@ 8 Conclusion
In this work, we revisit masked diffusion models, focusing on a flexible continuous-time formulation. Existing works in this area are not easily accessible to non-specialists and present ELBOs that are difficult to optimize, often resulting in performance that is not competitive with continuous diffusions and AR models. The framework we propose provides a very simple expression of the ELBO as a weighted integral of cross-entropy losses. Additionally, we propose a generalized masked diffusion formulation (GenMD4), where the masking schedule depends on the current state of the process, and derive its corresponding ELBO. On text data, our MD4 models outperform existing discrete and continuous diffusion models. For pixel-level image modeling, we significantly improve discrete diffusion results, outperforming similar-sized AR models and achieving comparable likelihoods to continuous diffusion models such as VDM. GenMD4 provides further improvements in terms of likelihoods over the state-independent case.
Although we have improved masked diffusion models, they still suffer from limitations. First, in some tasks such as text8, masked diffusions are not yet competitive with AR models. We conjecture that this is because AR models can better leverage model capacity since they only require learning one order. It would be interesting to develop better architectures for discrete diffusions. Moreover, GenMD4 is promising, but it can easily overfit to the dataset, making it less effective for zero-shot transfer compared to simpler versions. Additionally, inference with a state-dependent schedule is more challenging.
@@ H.3 Score parameterization
$\mathcal{L}_{\infty}=\int_{t(1)}^{1}\mathbb{E}_{q_{t|0}(k|x_{0})}\Big{[}\sum_{j\neq k}Q(t)_{jk}\Big{(}s_{\theta}(k)_{j}-\frac{q_{t|0}(j|x_{0})}{q_{t|0}(k|x_{0})}\log s_{\theta}(k)_{j}+\psi\Big{(}\frac{q_{t|0}(j|x_{0})}{q_{t|0}(k|x_{0})}\Big{)}\Big{)}\Big{]}\mathrm{d}t,$ (24)
@@
Note that only the result under the case $k=m$ is needed. This is because when $x_{t}$ is unmasked, at any time between 00 and $t$, the state must stay unchanged and remain $x_{0}$. As a result, $\mathrm{KL}({q(x_{t-\Delta t}|x_{t},x_{0})}\|{p_{\theta}(x_{t-\Delta t}|x_{t})})=0$ for $x_{t}\neq m$. From (15), we know $Q(t)_{jk}=\beta(t)(\delta_{mk}-\delta_{jk})$. Combining (25) and (24), we get
@@
This is the same as the conditional reverse transition rate used in Campbell et al. [49, Eq. (22)]—note that their time $t$ is reversed, and the rate matrix was therefore in the form $R^{x_{0}}(t)=\frac{1}{1-t}e_{m}(x_{0}-e_{m})^{\top}$.
@@ I.2 Training and gradient estimation
$\mathcal{L}^{(N)}_{\infty}=\int_{0}^{1}\Big{(}\frac{\alpha_{t}^{\prime}}{1-\alpha_{t}}\Big{)}^{\top}\mathbb{E}_{q(x_{t}|x_{0})}\Big{[}{\textstyle\sum}_{n:x_{t}^{(n)}=m}(x_{0}^{(n)}-\mu_{\theta}^{(n)}(x_{t},t)+x_{0}^{(n)}(x_{0}^{(n)})^{\top}\log\mu_{\theta}^{(n)}(x_{t},t))\Big{]}\mathrm{d}t.$ (32)
@@
$\mathcal{L}^{(N)}_{\infty}=-\int_{0}^{1}\frac{1}{t}w^{\top}\mathbb{E}_{q(x_{t}|x_{0})}\left[{\textstyle\sum}_{n:x_{t}^{(n)}=m}(x_{0}^{(n)}-\mu_{\theta}^{(n)}(x_{t},t)+x_{0}^{(n)}(x_{0}^{(n)})^{\top}\log\mu_{\theta}^{(n)}(x_{t},t))\right]\mathrm{d}t,$
@@
$-\frac{1}{2t}\left\{g(x_{t}^{1},x_{0})+g(x_{t}^{2},x_{0})+\left(\nabla_{w}\log q(x_{t}^{1}|x_{0})-\nabla_{w}\log q(x_{t}^{2}|x_{0})\right)\left[f(x_{t}^{1},x_{0})-f(x_{t}^{2},x_{0})\right]\right\}.$
@@
J.3 FineWeb-Edu
We kept the same training setup as the OpenWebText experiments. Our transformer models have the same number of layers, heads, and hidden dimensions as those of GPT-2 models. We use the same GPT-2 tokenizer.
For Hellaswag evaluation, we concatenate question with each answer option, tokenize the concatenated string, pad to the length of 1024. The padded token sequence gets fed to our MD4 model’s loss function for likelihood evaluation. We average 32 Monte Carlo samples to reduce variance. The answer with the highest likelihood estimate is the model’s prediction.
J.4 Images
@@
As explained in Sec. 4, we have observed that the cosine schedule leads to better sample quality so we used it to train a cheaper model for sample visualization. This model differs from the one that achieves best likelihood in that we used 8 residual blocks (in one side of the UNet structure) and a 20-layer diffusion transformer in the middle. All other configurations are kept the same.
@@
K.3 FID evaluation of MD4 trained on ImageNet 64$\times$64
We provide the FID numbers corresponding to Fig. 2 in Tab. 6.
Table 6: FID of 50k samples generated by MD4 trained on ImageNet 64$\times$ 64, corresponding to Fig. 2. Top three rows show results from an unconditional model, while the bottom row is from a model conditioned on class labels. Uniform discretization grid is used in Alg. 2 unless otherwise noted.
Method Timesteps $T$
64 128 256 512
Linear $\alpha_{t}$ 193.81 128.18 72.94 50.21
Linear $\alpha_{t}$, cosine grid 42.07 25.16 18.31 18.22
Cosine $\alpha_{t}$ 47.46 23.84 17.8 18.74
Cosine $\alpha_{t}$, class conditional 30.75 11.39 7.13 7.8
K.4 Additional unconditional generation from MD4 trained on ImageNet 64$\times$64
@@
K.5 Additional unconditional generation from MD4 trained on OpenWebText
@@
K.5.1 MD4-M unconditional sample 1: 1024 tokens
@@
K.5.2 MD4-M unconditional sample 2: 1024 tokens
@@
K.6 Conditional generation from MD4 trained on OpenWebText
@@
Figure 10: Conditionally generated text samples from MD4-M. Top: MD4-M trained with the linear schedule, sampled with a uniform grid; Middle: MD4-M trained with the linear schedule, sampled with the cosine grid; Bottom: MD4-M trained with the cosine schedule, sampled with a uniform grid. Context text shown in blue, model-generated text in black.
K.7 Effect of discretization on zero-shot perplexity